if the difference between the smallest and the largest of five consecutive odd number is t what is t-2
Answer by HyperBrain(694)

Let m-4, m-2, m, m+2, m+4 be the consecutive odd integers.
The largest is m+4 and the smallest is m-4.
(m+4)-(m-4)=8=t
Since t=8, then t-2=8-2=6
Power up,
HyperBrain!